Select any imaginary number (of the form "a + bi" where a and b are non-zero real numbers), and turn it into a real number by using addition, subtraction, multiplication, division, or exponentiation.
Answer by RAY100(1637)

The complex conjugate operation applies
.
(a+bi) (a-bi) = a^2 -abi +abi -b^2i^2 = a^2 +b^2
.
with i^2 = -1
.